Show that ,f defined on [a,b] is of bounded variation on [a,b] if and only if f can be expressed as the difference of two increasing functions


1
Expert's answer
2021-03-16T11:31:23-0400

Assume that "f \in V[a,b]" and let "v(x)=V^x_a(f), x \in (a,b]" and "v(a)=0". Then clearly "f(x) = v(x)-[v(x)-f(x)]". We will show "v(x)" and "v(x)-f(x)" are increasing.

"\forall" "x_1<x_2" : "v(x_2)-v(x_1)=V ^{x_2}_{x_1} (f) ≥ 0 \iff v(x_2) ≥ v(x_1)" so "v(x)" is increasing.

"\forall" "x_1<x_2" : "f(x_2)-f(x_1) \le |f(x_2)-f(x_1)| \le V ^{x_2}_{x_1} (f) =v(x_2)-v(x_1)" "\iff v(x_1)-f(x_1) \le v(x_2)-f(x_2)" and thus "v(x)-f(x)" is increasing.

Conversely, suppose that "f(x)=g(x)-h(x)" with "g" and "h" increasing. Since "h" is increasing, "-h" is decreasing and thus "f(x) = g(x)+(-h(x))" is the sum of two monotone functions. So theorem

"[" If "f : [a,b] \to \R" is monotone on "[a,b]" then "f \in V[a,b]" and "V^b_a(f)=|f(b)-f(a)|" "]" together with theorem "[" Let "f,g:[a,b] \to \R" be of bounded variation on "[a,b]". Then "(f+g)\in V[a,b]" and "V^b_a(f+g) \le V^b_a(f) + V^b_a(g)" "]" gives that "f \in V[a,b]".